OpenStudy (aglna):

Simply expression 3^3 X 3^3 / 3^4 will give meadal

OpenStudy (aglna):

Medal

OpenStudy (raffle_snaffle):

 Does this look right?

OpenStudy (aglna):

Yep

OpenStudy (raffle_snaffle):

so what is (3*3*3)*(3*3*3) = ?

OpenStudy (aglna):

36

OpenStudy (raffle_snaffle):

remember 3^3 = (3*3*3) <---- these are equal 3 is being raised to the third power meaning it needs to be multiplied by itself 3 times.